# On the interplay between the BFKL resummation and high-energy factorization in Mueller-Navelet dijet production

### v1

@@ Abstract
$\times$ $\delta(1-x/x_{g/q})\ \delta^{(2)}({\bf q}_{T}\mp{\bf l}_{T}-{\bf p}_{T}).$
Due to the lack of conformal symmetry in QCD, the complete basis of orthogonal eigenfunctions $H_{n,\gamma}$ of the BFKL kernel $K$ in Eq. (1) must be constructed order-by-order in $\alpha_{s}$, ensuring the RG-invariance of the Green’s function at NLL accuracy as Chirilli and Kovchegov (2013, 2014):
@@
$\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(n,\gamma)\,H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}).$
@@
$\chi^{(0)}(n,\gamma)$ $=$ $2\psi(1)-2\,{\rm Re}\,\psi\left(\frac{n}{2}+\gamma\right),$
$\chi^{(1)}(n,\gamma)$ $=$ $\frac{1}{2}\,\frac{\beta_{0}}{4\pi}\,\frac{\partial\chi^{(0)}(n,\gamma)}{\partial\gamma}+\frac{\delta(n,\gamma)}{4},$
the function $\delta(n,\gamma)$ is given by Eq. (27) in Ref. Kotikov and Lipatov (2000). Using the technique of Refs. Chirilli and Kovchegov (2013, 2014), the NLO eigenfunctions $H_{n,\gamma}$ (6) are constructed as:
$H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})=f^{(0)}_{n,\gamma}({\bf l}_{T})+\bar{\alpha}_{s}(\mu_{R}^{2})\,f^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}),$ (8)
@@
$f^{(0)}_{n,\gamma}({\bf l}_{T})$ $=$ $\frac{1}{\sqrt{\pi}}\,{\bf l}_{T}^{2(\gamma-1)}e^{in\phi_{l}},$
$f^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})$ $=$ $\sum\limits_{m=0}^{\infty}\,c_{m}(n,\gamma)\ln^{m}\left(\frac{{\bf l}_{T}^{2}}{\mu_{R}^{2}}\right)f^{(0)}_{n,\gamma}({\bf l}_{T}),$
with: $c_{0}=0$, $c_{1}(n,\gamma)=\partial_{\gamma}c_{2}(n,\gamma)$, $c_{2}(n,\gamma)=\beta_{0}/\left(8\pi\,\partial_{\gamma}\ln\chi^{(0)}(n,\gamma)\right)$, and $c_{m>2}=0$ at the NLO. The integral along the contour ${\rm Re}\,\gamma=1/2$ in Eq. (4) should be understood in the sense of the principal value treating the eigenfunctions as distributions due to the pole of the $c_{2}$ at $\gamma=1/2$.
Figure 3: The (a) characteristic function (7) at $\text{Re}\,\gamma=1/2$ as function of $\text{Im}\,\gamma$ for $n=0$ and $\mu_{R}=10$ GeV and (b) intercept dependence on $\mu_{R}$. Several options are shown: LO (dashed line), NLO (dot-dashed line), and CI NLO (solid line).
The Pomeron intercept characterizing the behavior of the Green’s function (4) at $Y\to\infty$ is defined at the saddle point $\gamma$ of the appropriate characteristic function (7): $\omega=\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(0,\gamma)$. At the LO, the latter one has only one saddle point at $\gamma=1/2$, which splits into two complex conjugate saddle points at the NLO, whose position depends on $\alpha_{s}$, see Fig. 3 (a). Our results for the characteristic function and intercept are consistent with ones presented in the recent Ref. Polizzi et al. (2025).
Figure 4: The plots of the function (10) calculated with the LL (dashed line) and CI NLL (solid line) Green’s function. In all plots except the ones showing the $\Delta\psi$-dependence, the function $g$ averaged over $\Delta\psi$ is plotted.
As it has been first observed in Ref. Salam (1998), the numerically largest NLO contribution to the characteristic function comes from (anti-)collinear poles, located at $\gamma=-n/2$ ($21+n/$ 2), which can be written as:
@@
with $\bar{\alpha}_{s}=\bar{\alpha}_{s}(\mu_{R}^{2})$, $d_{m}=d_{m}(n,\gamma)=m+\gamma+n/2$, and the coefficients $\kappa_{l,k}=\kappa_{l,k}(n)$ being defined from the pole structure of the fixed-order terms in Eq. (7). The effects of leading and sub-leading poles can be resummed Salam (1998); Sabio Vera (2005); Sabio Vera and Schwennsen (2007) to all orders in $\alpha_{s}$, the “all-poles” approximation of Refs. Sabio Vera (2005); Sabio Vera and Schwennsen (2007) for such a resummation has the form:
@@
The CI NLO characteristic function has only one saddle point preserving the LO definition of the intercept at $\gamma=1/2$, see Fig. 3 (solid line). Both, NLO and CI NLO intercepts, are approximately scale independent at relatively large $\mu_{R}$ typical for the MN dijet production and are much smaller than the LO intercept.
It is convenient, in particular for numerical interpolation purposes, to work with the function
$g\left({\bf l}_{T1},{\bf l}_{T2},Y\right)$ $=$ $\frac{|{\bf l}_{T1}||{\bf l}_{T2}|\,G\left({\bf l}_{T1},{\bf l}_{T2},Y\right)}{G_{\rm D}\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|,Y\right)},$ (10)
which is the exact Green’s function (4) divided by the diffusion approximation valid at $\bar{Y}\sim 1$, in LL approximation:
$G_{\rm D}\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|,Y\right)$ $=$ $\frac{1}{2\pi^{2}}\sqrt{\frac{\pi}{D\,\bar{Y}}}$
$\times$ $\exp\left[\omega_{0}\,\bar{Y}-\frac{1}{D\,\bar{Y}}\,\ln^{2}\left(\frac{|{\bf l}_{T1}|}{|{\bf l}_{T2}|}\right)\right],$
here $\omega_{0}=4\ln(2)$ and $D=14\zeta(3)$. The dependence of the function (10) on the momenta ${\bf l}_{T1,2}$ can be conveniently parameterized with the help of variables: $L=\ln\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|\right)$, $\Delta L=\ln\left(|\Delta{\bf l}_{T}|/\mu_{R}\right)$, where $\Delta{\bf l}_{T}={\bf l}_{T1}-{\bf l}_{T2}$, and $\Delta\psi$ being the azimuthal angle between ${\bf l}_{T1,2}$.
Several plots of the function (10) are shown in Fig. 4. At relatively small $\bar{Y}\lesssim 0.5$, the behavior of the function $g$ w.r.t. $L$ and $\Delta\psi$ approximates the initial condition of the Green’s function (4): $g\sim\delta(L)\,\delta(\Delta\psi)$. With rapidity increasing up to $\bar{Y}\sim 1.5$, the peak spreads and the LL $g$ function approaches unity due to the validity of the diffusion approximation in this region. The CI NLL Green’s function does not show any unphysical oscillations or negative values, as evident from Fig. 4. We also point out that the effect of CI on the Green’s function is numerically small, consistently with observations of Refs. Colferai et al. (2010); Ducloue et al. (2013); Ducloué et al. (2014) and the NLO approximation (7) is actually sufficient for the description of the data. The dependence on $\Delta L$ of the NLL $g$ is rather small, so we do not vary it in Fig. 4 setting $\Delta L=0$, i.e. $\mu_{R}=|\Delta{\bf l}_{T}|$. For $\bar{Y}\sim 1$, the NLL $g$-function is almost flat, but a few times lower than the LL one due to the smaller intercept value (Fig. 3 (b)). Thus, the full NLL Green’s function ($G$) grows slower with increasing of $\bar{Y}$ than the LL one.

### v2

@@ Appendix A End Matter materials
$\times$ $\delta(1-x/x_{g/q})\ \delta^{(2)}({\bf q}_{T}\mp{\bf l}_{T}-{\bf p}_{T})$
as in Eq. (3), but with $G$ replaced by $\delta^{(2)}\left({\bf l}_{T1}-{\bf l}_{T2}\right)$, which is the initial condition of the Green’s function in Eq. (4). This is the BFKL(0) HSC in Eq. (5), see Fig. 2 (b). Since the above-mentioned initial condition is included into the solution 4, we have to subtract the BFKL(0)-contribution from the HSC in Eq. (5) to avoid the double-counting between the HEF HSC and BFKL-improved HSC at $Y\sim 1/\bar{\alpha}_{s}(\mu_{R}^{2})$. After that we can add to Eq. (5) the BFKL-resummed contribution (3) with the full Green’s function in Eq. (4), drawn in Fig. 2 (c).
Due to the lack of conformal symmetry in QCD, the complete basis of orthogonal eigenfunctions $H_{n,\gamma}$ of the BFKL kernel $K$ in Eq. (1) must be constructed order-by-order in $\bar{\alpha}_{s}$, ensuring the RG-invariance of the Green’s function at the NLL accuracy as Chirilli and Kovchegov (2013, 2014):
@@
$\hskip-56.9055pt\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(n,\gamma)\,H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}).$
@@
$\chi^{(0)}(n,\gamma)$ $=$ $2\psi(1)-2\,{\rm Re}\,\psi\left(\gamma+\frac{n}{2}\right),$
$\chi^{(1)}(n,\gamma)$ $=$ $\frac{\beta_{0}}{8\pi}\frac{\partial\chi^{(0)}(n,\gamma)}{\partial\gamma}+\frac{\delta(n,\gamma)}{4},$
the function $\delta(n,\gamma)$ is given by Eq. (27) in Ref. Kotikov and Lipatov (2000). An important property of $\chi(n,\gamma)$ is it’s symmetry w.r.t. $\gamma\leftrightarrow 1-\gamma$, which is necessary for the existence of the decomposition in Eq. (4). Using the technique of Refs. Chirilli and Kovchegov (2013, 2014), the NLO eigenfunctions $H_{n,\gamma}$ (6) are constructed as:
$H_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})=H^{(0)}_{n,\gamma}({\bf l}_{T})+\bar{\alpha}_{s}(\mu_{R}^{2})\,H^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2}),$ (8)
@@
$H^{(0)}_{n,\gamma}({\bf l}_{T})$ $=$ $\frac{1}{\sqrt{\pi}}\,{\bf l}_{T}^{2(\gamma-1)}e^{in\phi_{l}},$
$H^{(1)}_{n,\gamma}({\bf l}_{T},\mu_{R}^{2})$ $=$ $\sum\limits_{m=0}^{\infty}\,c_{m}(n,\gamma)\ln^{m}\left(\frac{{\bf l}_{T}^{2}}{\mu_{R}^{2}}\right)H^{(0)}_{n,\gamma}({\bf l}_{T}),$
with: $c_{0}=0$, $c_{1}(n,\gamma)=\partial_{\gamma}c_{2}(n,\gamma)$, $c_{2}(n,\gamma)=\beta_{0}/\left(8\pi\,\partial_{\gamma}\ln\chi^{(0)}(n,\gamma)\right)$, and $c_{m\geq 3}=0$ at the NLO. The integral along the contour ${\rm Re}\,\gamma=1/2$ in Eq. (4) should be understood in the sense of the principal value treating the eigenfunctions as distributions due to the pole of the $c_{2}$ at $\gamma=1/2$.
Figure 3: The (a) characteristic function (7) at $\text{Re}\,\gamma=1/2$ as function of $\text{Im}\,\gamma$ for $n=0$ and $\bar{\alpha}_{s}=0.2$ and (b) intercept dependence on $\mu_{R}$. Several options are shown: LO (dashed line), NLO (dot-dashed line), and CI NLO (solid line).
Figure 4: The plots of the LL and NLL Green’s function (4). In all plots except the ones showing the $\Delta\psi$-dependence, the function is averaged over $\Delta\psi$ is plotted. The value $\mu_{R}=10$ GeV is fixed, except the case of the dependence on it.
As it has been first observed in Ref. Salam (1998), the numerically largest NLO contribution to the characteristic function comes from (anti-)collinear poles, located at $\gamma=-n/2$ ($1+n/2$), which can be written as:
@@
where $\bar{\alpha}_{s}=\bar{\alpha}_{s}(\mu_{R}^{2})$, $d_{m}=d_{m}(n,\gamma)=m+\gamma+n/2$, and the coefficients $\kappa_{l,k}=\kappa_{l,k}(n)$ being defined from the pole structure of the fixed-order terms in Eq. (7). The effects of leading and sub-leading poles can be resummed Salam (1998); Sabio Vera (2005); Sabio Vera and Schwennsen (2007) to all orders in $\alpha_{s}$, the “all-poles” approximation of Refs. Sabio Vera (2005); Sabio Vera and Schwennsen (2007) for such a resummation has the form:
@@
The Pomeron intercept characterizing the behavior of the Green’s function (4) at $Y\to\infty$ is defined at the saddle point $\gamma$ of the appropriate characteristic function (7): $\omega=\bar{\alpha}_{s}(\mu_{R}^{2})\,\chi(0,\gamma)$. At the LO, the latter one has only one saddle point at $\gamma=1/2$, which splits into two complex conjugate points at the NLO, whose position depends on $\bar{\alpha}_{s}$, see Fig. 3. Our results for the characteristic function and intercept are consistent with ones presented in the recent Ref. Polizzi et al. (2025). The CI NLO characteristic function has only one saddle point preserving the LO definition of the intercept, see Fig. 3 (solid line). Both, NLO and CI NLO intercepts, feature a reduced scale-dependence at relatively large $\mu_{R}$ typical for the MN dijet production in comparison with the LO intercept and are much smaller than the LO intercept.
Several plots of the LL and NLL Green’s function (4) are shown in Fig. 4. At relatively small $\bar{Y}\sim 0.5$, the behaviour of the function w.r.t. $L=\ln\left(|{\bf l}_{T1}|/|{\bf l}_{T2}|\right)$ (Fig. 4 (a)–(c)) and $\Delta\psi$ (Fig. 4 (d)–(f)) being the azimuthal angle between ${\bf l}_{T1,2}$, approximates the initial condition (4) and with rapidity increasing up to $\bar{Y}\sim 1.5$, the peak spreads. The CI NLL function does not show any unphysical oscillations or negative values, as evident from Fig. 4. The effect of the CI is numerically small, consistently with observations of Refs. Colferai et al. (2010); Ducloue et al. (2013); Ducloué et al. (2014). The NLL function grows slower with the increase of $\bar{Y}$ than the LL one due to the smaller intercept value (Fig. 4 (g)–(i)). Finally, we verify the RG-invariance of the Green’s function at the NLO accuracy in the region $\bar{Y}\sim 1$ (Fig. 4 (j)–(l)).
@@
We have checked the consistency of the solution (4) with the running-coupling solution proposed in Ref. Grabovsky (2013) and found the agreement up to NLL, as one would expect. We will present the details of this comparison elsewhere due to the lack of space.